OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 3x2 + 6y2 = 18 @amistre64 @mathstudent55 help me please

OpenStudy (anonymous):

Medal :)

OpenStudy (anonymous):

I'm Guessing it's this Center: (0, 0); Vertices: point negative square root of six comma zero and point square root of six comma zero; Foci: Ordered pair negative square root three comma zero and ordered pair square root three comma zero

OpenStudy (amistre64):

the center should be obvious

OpenStudy (anonymous):

Center: (0, 0); Vertices: point negative square root of six comma zero and point square root of six comma zero; Foci: Ordered pair negative square root three comma zero and ordered pair square root three comma zero I'm guessing its this

OpenStudy (amistre64):

since the center is 0,0 then the vertexes will simply be the solutions for x=0, solve for y .... and y=0, solve for x

OpenStudy (amistre64):

foci are trickier, but doable rewriting the equation into an =1 format we can read the bottoms under the x and y parts 3x2 + 6y2 = 18 3x2 6y2 18 ---- + ---- = ---- 18 18 18 x2 y2 ----- + ----- = 1 18/3 18/6 x2/6 + y2/3 = 1 now the foci are along along the x axis, since under x is larger, the distance is detemrined by some malformed pythagorean thrm: f^2 + x^2 = 6^2 will determine f for us

OpenStudy (amistre64):

last line is a typo f^2 + 3^2 = 6^2, solves for f

OpenStudy (amistre64):

youve guessed at the foci, but we have 4 verts ... 2 you named correctly, and you called the other 2 the foci :)

OpenStudy (amistre64):